# Antiferromagnetic pseudogap in the two-dimensional Hubbard model deep in the renormalized classical regime

### v1

@@ IV Conclusion
The results for two-particle properties are in semi-quantitative agreement with DiagMC and $D\Gamma A$ results in the region where the latter converge. We were able to go to much lower temperatures than most other theoretical approaches and to calculate the spin correlation length $\xi$ and spin-susceptibility $\chi_{sp}$ in the thermodynamic limit for temperatures with extremely large correlation lenghts $\xi\leq 10^{6}$. This requires special numerical approaches, discussed in Appendix A.
@@
Our three main conclusions are that (a) For the single-particle properties, the TPSC+Classical approach allows us to describe quantitatively DiagMC data in both the metallic and the pseudogap regime. (b) The self-energy results at very low temperatures lead to estimates for the zero temperature gap that is smaller than the renormalized mean-field result (for $U=2,n=1$) by about 20%, due to quantum fluctuations. (c) Analytic continuation of the Matsubara frequency results for the single particle properties using Padé approximants show that the pseudogap in the spectral function opens up significantly earlier than what was determined previously using Matsubara frequency data only and that the criterion $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$ is a sufficient but not a necessary condition for the pseudogap.
The main puzzle that remains to be solved is the following. In the pseudogap regime, the spectral function at $T\rightarrow 0$ has two peaks at $\omega\pm\Delta$ and an incoherent background due to quantum fluctuations. For $|\omega|>\Delta$ the presence of the incoherent background is consistent with previous results from the $NL\sigma M$ (non-linear sigma model) of the antiferromagnetic fluctuations [borejsza2003antiferromagnetism]. Our theory also predicts a very small but finite spectral weight for $|\omega|<\Delta$ and $T\rightarrow 0$ due to quantum fluctuations. This implies a discontinuity between the zero-temperature limit of the spectral weight in the pseudogap regime and the spectral weight in the zero-temperature long-range antiferromagnetic state. The small spectral weight in the antiferromagnetic pseudogap is likely an artifact of the one loop approximation for the self-energy. Designing a theory that completely bridges the pseudogap regime and the zero-temperature antiferromagnetic state is an open problem.
@@ A.2 Thermodynamic Limit
The 2D integral of this expression has a quasi-singularity at $\mathbf{q}=\mathbf{Q}$ and thus equally-spaced integration routines do not work. Since the quasi-singularity has circular symmetry, we convert the integral over $\mathbf{q}$ to cylindrical coordinates relative to $\mathbf{q}=\mathbf{Q}$:
@@
The integrand in this integral does not have a quasi-singularity at the lower limit and it can be evaluated using standard integration routines. Note, that the number of function evaluations increases with $\xi$ but only logarithmically, which effectively solves the problem.
@@
However, if one is interested in the Matsubara self-energy at very low temperature to find the zero temperature gap, one can suppress the infinite-lattice step, but allow accurate finite mesh calculations of the Matsubara self-energy, of double-occupancy and of other physical properties whose dependence on finite system size is relatively weak.

### v2

@@ IV Conclusion
The results for two-particle properties are in semiquantitative agreement with DiagMC and $D\Gamma A$ results in the region where the latter converge. We were able to go to much lower temperatures than most other theoretical approaches and to calculate the spin correlation length $\xi$ and spin-susceptibility $\chi_{sp}$ in the thermodynamic limit for temperatures with extremely large correlation lenghts $\xi\leq 10^{6}$. This requires special numerical approaches, discussed in Appendix A.
@@
Our three main conclusions are that (a) For the single-particle properties, the TPSC+Classical approach allows us to describe quantitatively DiagMC data in both the metallic and the pseudogap regime. (b) The self-energy results at very low temperatures lead to estimates for the zero-temperature gap that is smaller than the renormalized mean-field result (for $U=2,n=1$) by about 20%, due to quantum fluctuations. (c) Analytic continuation of the Matsubara frequency results for the single particle properties using Padé approximants show that the pseudogap in the spectral function opens up significantly earlier than what was determined previously using Matsubara frequency data only and that the criterion $|\Sigma(\mathbf{k}_{F},ik_{0})|>|\Sigma(\mathbf{k}_{F},ik_{1})|$ is a sufficient but not a necessary condition for the pseudogap.
The main puzzle that remains to be solved is the following. In the pseudogap regime, the spectral function at $T\rightarrow 0$ has two peaks at $\omega\pm\Delta$ and an incoherent background due to quantum fluctuations. For $|\omega|>\Delta$ the presence of the incoherent background is consistent with previous results from the $NL\sigma M$ (nonlinear sigma model) of the antiferromagnetic fluctuations [66]. Our theory also predicts a very small but finite spectral weight for $|\omega|<\Delta$ and $T\rightarrow 0$ due to quantum fluctuations. This implies a discontinuity between the zero-temperature limit of the spectral weight in the pseudogap regime and the spectral weight in the zero-temperature long-range antiferromagnetic state. The small spectral weight in the antiferromagnetic pseudogap is likely an artifact of the one loop approximation for the self-energy. Designing a theory that completely bridges the pseudogap regime and the zero-temperature antiferromagnetic state is an open problem.
@@ A.2 Thermodynamic Limit
The 2D integral of this expression has a quasisingularity at $\mathbf{q}=\mathbf{Q}$ and thus equally-spaced integration routines do not work. Since the quasisingularity has circular symmetry, we convert the integral over $\mathbf{q}$ to cylindrical coordinates relative to $\mathbf{q}=\mathbf{Q}$:
@@
The integrand in this integral does not have a quasisingularity at the lower limit and it can be evaluated using standard integration routines. Note, that the number of function evaluations increases with $\xi$ but only logarithmically, which effectively solves the problem.
@@
However, if one is interested in the Matsubara self-energy at very low temperature to find the zero-temperature gap, one can suppress the infinite-lattice step, but allow accurate finite mesh calculations of the Matsubara self-energy, of double-occupancy and of other physical properties whose dependence on finite system size is relatively weak.